OpenStudy (anonymous):

express as a single fraction the exact value of sin75

OpenStudy (anonymous):

\[\huge \sin75^0 = \sin (45^0+ 30^0)\] \[\huge= \sin 45^0 \cos 30^0 + \cos45^0 \sin 30^0\] \[\huge= \frac{1}{ \sqrt2} \frac{\sqrt3}{2} + \frac{1}{ \sqrt2} \frac{1}{2} \] \[\huge= \frac{\sqrt3}{2\sqrt2} + \frac{1}{ 2\sqrt2} = \frac{\sqrt3 +1}{2\sqrt2}\] is the required answer.

OpenStudy (anonymous):

@awsome500

OpenStudy (anonymous):

Since 75= 45+30

OpenStudy (anonymous):

so i used 45 + 30 in place of 75 and also we don't know the direct value of 75 degrees so we need to split 75 as 45+ 30

OpenStudy (anonymous):

@awsome500 any further query

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

@awsome500 Medal please...

OpenStudy (anonymous):

cos 45 = \[\frac{ \sqrt{2} }{ 2}\]

OpenStudy (anonymous):

sin75 will the answer be \[\frac{ \sqrt{6}+\sqrt{2} }{ 4 }\] @mathmale

OpenStudy (mathmale):

Without going through every detail of the other person's attempts to show you the way, I'll say his approach does look very appropriate. Why are you in doubt over his answer? I'd suggest you review his work and ask questions about anything not clear to you. I'd be delighted to respond in that case. I'd rather not be asked to go through someone else's entire solution of a problem and judge whether it's correct or not.

OpenStudy (whpalmer4):

@awsome500 both of those answers are equivalent; the one you propose is simply the one @dpasingh gave after being rationalized.

OpenStudy (anonymous):

thank you i understand

OpenStudy (whpalmer4):

\[\frac{1+\sqrt{3}}{2\sqrt{2}} = \frac{1+\sqrt{3}}{2\sqrt{2}} *\frac{\sqrt{2}}{\sqrt{2}} = \frac{\sqrt{2}(1+\sqrt{3})}{2\sqrt{2}\sqrt{2}} = \frac{\sqrt{2}+\sqrt{6}}{4}\]

OpenStudy (whpalmer4):

Unless the problem specifically requests that you rationalize the denominator, either one should be acceptable.

OpenStudy (anonymous):

@mathmale please help me on one more question express \[\frac{ \cot \chi \sin \chi }{ \sec }\] as a single trigonometric function in simplest form for all values of x for which it is defined

OpenStudy (mathmale):

Both the secant and the cotangent can be re-written in terms of sine and cosine. Would you please do that now, before considering any other action?

OpenStudy (anonymous):

okay cot will be cos over sin and sec will be 1 over cos

OpenStudy (mathmale):

Good! Now let's put that all together. Here's what we now have:\[\frac{ \cos x }{ \sin x}\frac{ \sin x }{ 1 } ~divided~by~\frac{ 1 }{ \cos x }\]

OpenStudy (mathmale):

Try writing that out on paper. Cancel wherever possible.

OpenStudy (anonymous):

so we cancel sin which will leave us with cos over 1 divided by 1over cos

OpenStudy (anonymous):

doesnt cos get canceled out to

OpenStudy (mathmale):

I'd like to see what you have done; then I'd be in a better position to answer that. Can you upload an image? If not, type out what you have done.

OpenStudy (anonymous):



OpenStudy (solomonzelman):

Yeah, so you canceled sin(x) and now you have \(\Huge\color{blue}{ \cal \frac{Cos(x)}{1}=\frac{Cos(x)}{1} }\) well, you reached the point a=a .

OpenStudy (anonymous):

but sec = 1over cos @SolomonZelman

OpenStudy (solomonzelman):

I didn't see the actual thing, you drew cos x /1 on the right side.. OK. So you have \(\Huge\color{blue}{ \cal \frac{Cos(x)}{1}=\frac{1}{Cos(x)} }\) you would cross multiply. Tell me what you get ?

OpenStudy (solomonzelman):

Or, you can just use a common sense, knowing that Cos(x) can't equal anything else, other than 1.

OpenStudy (solomonzelman):

If you were proving/verifying, then both sides aren't equal. because they are not equal for all values of x.

OpenStudy (anonymous):

did u flip the recpical of sec which was 1 over cos and it became cos over 1

OpenStudy (mathmale):